For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation?
−x + 2y = 9
4x + 5y = 9

Found 4 solutions by lynnlo, richwmiller, ikleyn, MathTherapy:
Answer by lynnlo(4176)
Answer by richwmiller(17219)

add x
subtract 9
2y-9

Answer by ikleyn(53742)

.
For the following system, if you isolated x in the first equation to use the Substitution Method,
what expression would you substitute into the second equation?
-x + 2y = 9
4x + 5y = 9
~~~~~~~~~~~~~~~~~~~~~~~~

I isolate 'x' in the first equation by moving 'x' to the right side (with changing its sign)
and moving '9' to the left side (with changing its sign)

    2y - 9 = x.


Then precisely this expression I will substitute into the second equation, replacing 'x' there
with 2y-9.

Thus I answered your question.



Answer by MathTherapy(10801)

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression
would you substitute into the second equation?

−x + 2y = 9
4x + 5y = 9

−  x + 2y = 9

+ x - 9 - 9 + x ----- Adding "x" to, and subtracting 9 from both sides
 2y - 9 = x <=== 1st equation solved for "x"